3. Sketch the level curves of the function f (, y) = y2 + 2x + y2 - 64 corresponding to the z-values 6 and 6.
Answer #1

Z G + 2n +-64 -6 = 22 2 20 (42t1) (5-61+9)+6-1-9 =0 (-3 22 This an circle oith a dus centre (-1,3) for be Come pugtin sile 42(x+1)^2+(y-3)^2=4 Implicit (X+1)^2+(y-3)^2=16 Implicit 8 6 4 -2 -6 |-2 -4